# L4P: Low-Level 4D Vision Perception Unified

### v1

@@ 4.4 Sparse 2D/3D Track Estimation
We evaluate on TAPVid-3D [31], a benchmark containing around 2.1M long-range 3D point trajectories from over 4000 real-world videos, covering a variety of objects, camera and object motion patterns, and indoor and outdoor environments. It consists of three datasets: Aria [39], DriveTrack [51], and PStudio [24]. It introduced several baselines by combining SOTA 2D point tracking approaches, such as TAPIR [8], BootsTAPIR [9], and CoTracker [26], with depth solutions like ZoeDepth [2], a monocular depth estimation approach, and COLMAP [45, 47], a structure-from-motion pipeline. The top performing approach on the benchmark is SpaTracker [68].
The benchmark evaluates both 3D and 2D tracking approaches, and uses metrics that measure the ability to predict point visibility using an occlusion accuracy metric (OA), the accuracy of predicted trajectories in the visible regions (APD), and joint occlusion and geometric estimation (AJ). To resolve the scale ambiguity in depth estimation, the benchmark uses global median scaling by computing the median of the depth ratios between the estimated and ground-truth 3D tracks over all the points and frames in a video. We use the full_eval split evaluation numbers provided in the TAPVid-3D benchmark for comparing approaches.
On 3D tracking, we outperform previous approaches on average across all the metrics (Table 4). Among feedforward approaches, we perform better on all the datasets. Approaches that combine 2D track estimation with COLMAP perform better on the DriveTrack [51] dataset. This could be due to a relatively large bias of tracking mostly static vehicles, where COLMAP gives much more accurate depth. Such COLMAP-based baselines, however, perform poorly on Aria [39] and PStudio [24], which are mostly dynamic. We show qualitative evaluation against the SOTA SpaTracker approach in Figure 7.
On 2D tracking, we are slightly behind the SOTA 2D tracking approaches (Table 3). Our approach becomes more competitive and performs better than TAPIR on average when we fine-tune our model only for the 2D tracking task. We believe our reduced performance on 2D tracking comes from working at lower image resolution, $224\times 224$ for us as compared to $384\times 512$ for CoTracker and $256\times 256$ for others, and a lack of task-specific tricks, like tracking multiple points together (CoTracker) or assuming access to all frames in the video and performing a global track-refinement (TAPIR and BootsTAPIR), both of which could also benefit our tracking head. We also ablate our online tracking approach on both 2D and 3D tracking benchmarks, and show improved performance due to the use of memory mechanism when tracking points from one window to next. Overall, we attribute our strong performance to our unified approach and carefully designed sparse head.
@@ 4.5 Ablations
To understand the contribution of different components of our approach, we perform an ablation study for depth, flow, 2D and 3D point tracking, as shown in Table 5. For each of these tasks, we report average over the datasets not used in our training: for depth we use datasets in Table 1, for optical flow we use the Spring dataset, and for tracking we use the minival split from the TAPVid-3D [31] benchmark. Our main contribution is to show how to leverage the priors of a pretrained VideoMAE for multiple dense and sparse low-level 4D perception tasks at once. To show the usefulness of our end-to-end fine-tuning strategy, we compare against a pretrained and frozen VideoMAE, where we only fine-tune the task-specific heads. Table 5 shows that our fine-tuned VideoMAE (row 3) produces better results than the pretrained and frozen VideoMAE across all tasks (row 2). A version trained end-to-end from scratch results in worse performance (row 1), which shows that our system leverages the pretraining of the VideoMAE. Finally, by adding the proposed memory mechanism for the tracking head and using our two-stage training process, we obtain improvements in both 2D and 3D tracking tasks, while maintaining the performance on other tasks.
@@
From scratch 0.259 / 0.594 0.246 / 96.2 16.6 1.3
VideoMAE frozen 0.137 / 0.841 0.120 / 98.2 29.3 3.3
Ours (w/o Mem) 0.120 / 0.876 0.100 / 98.5 41.1 8.7
Ours 0.120 / 0.876 0.100 / 98.5 50.2 10.8
Table 5: Ablation study. Training using pre-trained VideoMAE performs better than training from scratch (row 3 vs. 1), which shows our approach leverages VideoMAE priors. Our approach performs better than using a frozen VideoMAE and only fine-tuning the heads (row 3 vs. 2), which shows end-to-end fine-tuning helps. Adding memory mechanism and two-stage training strategy improves tracking performance while maintaining performance on other tasks (row 4 vs. 3).
4.6 Additional Task: Motion-based Segmentation
We use the motion-based segmentation task to show one way to add a new task to our network. We do this simply by freezing our trained video encoder and fine-tuning our proposed dense head for this task. We generate the ground-truth annotations for training and evaluation by using video datasets that provide camera, depth and 3D-motion information. For training, we use the Kubric [16] dataset and fine-tune using binary cross entropy loss. For evaluation, we use the Virtual KITTI (VKITTI) [5] and Spring [37] datasets. We compare against RigidMask (RM) [72], a SOTA two-frame rigid-motion segmentation approach that combines dynamic motion signals from flow, optical-expansion [71] and depth. It is trained on the SceneFlowDatasets [36]; however, they also train a version for driving scenarios (RM-Drive). To evaluate, we report foreground IoU (higher is better) on VKITTI and Spring.
@@ 4.6 Additional Task: Motion-based Segmentation
RM 32.6 16.5
RM-Drive 35.4 8.5
Ours 46.7 23.7
On both datasets, our video-based approach achieves better performance. Note that while fine-tuning on driving scenes allows RigidMask (RM-Drive) to reduce the gap slightly on VKITTI, it significantly hurts performance on Spring, highlighting the benefit of our model’s generalization ability. As shown in Figure 8, for both the indoor scenarios with human-object interactions and the outdoor driving scenarios, our approach performs better and can detect small motions (see more comparisons in Supplementary). Freezing the video encoder and fine-tuning a task-specific head is the simplest way to add a new task that does not affect the performance of other tasks we train for. Better strategies may exist that allow for some fine-tuning of the video encoder without affecting the performance of other tasks, though the investigation is outside the scope of this paper.
Figure 8: Qualitative results of motion-based segmentation. Samples are chosen from the TAPVid-3D benchmark. Across various scenarios, ours show advantages in small motions, boundary accuracy, as well as temporal consistency (see Supp.). The GT masks overlaid on images are only provided to identify qualitatively which objects are moving, but are not pixel-accurate.
@@ 5 Conclusions
We present a unified framework to solve multiple low-level 4D vision perception tasks, both dense and sparse. We achieve this by adopting a strong pre-trained video masked auto-encoder and design lightweight task heads to harness its representation power. Our simple yet versatile designs for task heads allows for effortless and generalizable adaptation to multiple 4D vision perception tasks.
@@ 6 Acknowledgements
In this supplementary document, we provide comparisons for the inference time, and discuss additional implementation details regarding architecture, dataset and training. Please refer to our project webpage for a high-level overview and video results and comparisons.
@@ 1 Inference time
We compare inference time for several approaches we compared against in our paper in Table 1. For a video clip of size $16\times 224\times 224$, our inference runs in around 300ms on NVIDIA A6000 to generate output for all the tasks we address in our paper. While our approach does not provide latency low-enough for real-time applications, its inference time of 19ms per-frame makes it at least comparable speed-wise to methods specialized for each task. Our ability to solve multiple perception tasks under a reasonable computational envelope could be potentially useful for many applications like robotics, autonomous vehicles, etc.
@@ 2 Training datasets
Our video encoder [58] has been pre-trained on 1.35M video clips across various data sources using masked auto-encoding. To fine-tune our model, we use a limited number of synthetic datasets covering varying a range of annotations, and rely on the priors from the video encoder for generalization.
Kubric [16]. This synthetic dataset has multi-object interactions with many random objects. We use it to generate annotations for depth, flow, motion-based segmentation, and 2D/3D tracking. Each video is 24 frames long, and we use a total of 15k videos from the movi-e and movi-f subsets of the data. Kubric provides meta-data for object and camera trajectories, which could be used to generate 3D tracks. We follow official guidelines and generate the annotations for 3D tracking by sampling around 8-12k tracks in each video. For motion-based segmentation, we use the camera-pose and the 3D track information to detect which 3D tracks come from dynamic vs. static objects, which we then combine with provided rigid-object segmentations to generate annotations for motion-based segmentation.
PointOdyssey [77]. We use this synthetic dataset for depth and 2D/3D tracking annotations. The dataset consists of 159 videos, averaging 2k frames long, with human and object interactions under different lighting and atmospheric conditions. We sample smaller video clips from the long videos to form our dataset.
DynamicReplica [25]. We use this synthetic dataset for depth, flow and 2D/3D tracking annotations. The dataset consists of 524 videos with humans and animals in motion, and we sample smaller video clips to form our dataset. Since this dataset has higher fps videos, we sample videos with strides of 1, 2 and 4.
TartanAir [62]. Finally, to increase the scene-level data distribution we use TartanAir to generate annotations for flow and depth. The data is collected in photo-realistic simulation environments, with both indoor and outdoor scenes, in the presence of various light and weather conditions. We sample smaller video clips from this dataset.
@@
Table 1: Inference time. We compare our per-frame inference time with several task-specific approaches and show at least comparable speed-wise to methods specialized for each task. SpaTracker inference time is measured without the depth estimation.
@@ 3 Architecture
Our video encoder [58] processes video-clips of size $16\times 224\times 224$. It uses a patch-size of $2\times 14\times 14$, which results in $P=2048$ video tokens, and an embedding dimension of $C=1408$. It has 40 encoder blocks, and we use output from blocks $[14,21,28,36]$ for DPT heads for dense tasks, while our sparse-head uses features from the last block. We adapt DPT head [41] without any modifications, and only replace 2D-convolutions with 3D. For the sparse head, we adapt two-way attention from SAM [28] without any modifications and following SAM we keep this head lightweight by using only two two-way attention layers. We replace the 2D convolutions in the original mask-decoder of SAM with 3D convolutions. All our input/output tokens are of dimension $C=1408$ and use learnable embeddings. We use 3D positional encoding for the point query token and the video tokens for the two-way attention stage.
@@ 4 Training
We train in two-stages. In the first stage we train all the parameters of our model for depth, flow and 3D tracking on a single window of $T=16$ frames. During this stage, for each video we estimate all $T$ frames for dense tasks and for 3D tracking we construct a batch of 80 tracks and select point queries randomly across the visible parts of tracks. In the second stage, we further fine-tune our model using unrolled-window training. To train this stage efficiently, we only train for the 2D and 3D tracking tasks and freeze all the parameters, except the last three layers (37-39) of the video encoder and the sparse head. We train on videos of length $S=40$, windows of size 16 frames and a stride of 8, which results in 4 overlapping windows, and we construct a batch of 48 tracks during this stage, but generate point queries only in the first 20 frames to force the network to learn long-range tracking. We use AdamW optimizer with a maximum learning rate of $5\times 10^{-5}$ and a cosine annealing strategy. Both stages use a batch size of 8, are trained on a single 8-GPU NVIDIA A100 node for 100k iterations and take 1 day and 2 days respectively to train.

### v2

@@ 5.3 Sparse 2D/3D Track Estimation
We evaluate on TAPVid-3D [32], a benchmark containing around 2.1M long-range 3D point trajectories over three datasets: Aria [39], DriveTrack [53], and PStudio [25]. It introduced several baselines by combining SOTA 2D point tracking approaches, such as TAPIR [8], BootsTAPIR [9], and CoTracker [27], with depth solutions like ZoeDepth [2] and COLMAP [46, 48]. The top performing approach on the benchmark is SpaTracker [74].
The benchmark evaluates both 3D and 2D tracking approaches, and uses metrics that measure the ability to predict point visibility using an occlusion accuracy metric (OA), the accuracy of predicted trajectories in the visible regions (APD), and joint occlusion and geometric estimation (AJ). To resolve the scale ambiguity in depth estimation, the benchmark uses global median scaling by computing the median of the depth ratios between the estimated and ground-truth 3D tracks over all the points and frames in a video. We use the full_eval split evaluation numbers provided in the TAPVid-3D benchmark for the comparison.
On 3D tracking, we outperform previous approaches on average across all the metrics (Table 4). Approaches that combine 2D track estimation with COLMAP perform better on the DriveTrack [53] dataset. A reason could be that many query points are on static vehicles, for which COLMAP gives accurate depth. Such COLMAP-based baselines, however, perform poorly on Aria [39] and PStudio [25], which are mostly dynamic. We show a comparison with SpaTracker (SOTA) in Figure 6.
On 2D tracking, we are slightly behind the SOTA 2D tracking approaches (Table 3). Our approach becomes more competitive when we fine-tune our model only for the 2D tracking task. We believe our reduced performance on 2D tracking comes from working at lower image resolution, $224\times 224$ for us as compared to $384\times 512$ for CoTracker, and $256\times 256$ for others, and a lack of task-specific tricks, like tracking multiple points together (CoTracker), or assuming access to all frames in the video and performing a global track-refinement (TAPIR and BootsTAPIR), both of which could also benefit our tracking head.
@@ 5.4 Ablations
To understand the contribution of different components of our approach, we perform an ablation study for depth, flow and 3D point tracking, as shown in Table 5. For each of these tasks, we report average numbers over the datasets not used in our training: for depth we use datasets in Table 1, for optical flow we use the Spring dataset, and for tracking we use the minival split from the TAPVid-3D [32] benchmark. Our main contribution is a method that can leverage the priors of a pre-trained VideoMAE for dense and sparse low-level perception tasks. To show the usefulness of our end-to-end fine-tuning strategy, we compare against a pre-trained and frozen VideoMAE, where we only fine-tune the task-specific heads. Table 5 shows that our fine-tuned VideoMAE (row 4) produces better results than the pre-trained and frozen VideoMAE across all tasks (row 2). A version trained end-to-end from scratch results in worse performance (row 1), which shows that our system successfully builds on the priors of the pre-trained VideoMAE. Training our system for all tasks jointly (row 4) performs roughly on par as training for each task individually (row 3), which shows effectiveness of our training strategy to leverage VideoMAE priors for multiple tasks at once. Finally, by adding the proposed memory mechanism for tracking and using our multi-stage training process for online tracking, we obtain improvements for point tracking, while maintaining the performance on other tasks (row 5 vs. 4).
@@
L4P (from scratch, w/o mem) 0.274 / 0.560 0.319 / 95.0 15.9 / 1.2
L4P (VideoMAE frozen, w/o mem) 0.140 / 0.830 0.114 / 98.3 33.3 / 4.2
L4P (task-specific, w/o mem) 0.108 / 0.894 0.084 / 98.8 39.4 / 8.4
L4P (w/o mem) 0.103 / 0.895 0.092 / 98.7 40.4 / 8.3
L4P 0.103 / 0.895 0.092 / 98.7 49.0 / 10.7
Table 5: Ablation study. See Section 5.4 for an analysis.
5.5 Additional Tasks
L4P can be extended to new tasks, without degrading the performance of the original tasks. We do this by freezing our video encoder trained for the original tasks, and training a task-specific head. Here we provide two examples.
Motion-based segmentation. We train on the Kubric [16], and evaluate on the Virtual KITTI (VKITTI) [5] and Spring [37]. We compare against RigidMask (RM) [77], a SOTA two-frame approach. It is trained on the SceneFlowDatasets [36]; however, they also train a version for driving scenarios (RM-Drive). To evaluate, we report foreground IoU (higher is better).
@@ 5.5 Additional Tasks
RM $32.4$ $17.0$
RM-Drive $36.5$ $8.4$
L4P $\mathbf{56.0}$ $\mathbf{21.5}$
On both datasets, our video-based approach performs better. Note that while fine-tuning on driving scenes helps RigidMask (RM-Drive) on VKITTI, it significantly hurts performance on Spring, highlighting the benefit of our model’s generalization ability. As shown in Figure 5 (third row), our approach performs better and can detect small motions. See more comparisons in the Supplementary.
Camera pose estimation. We follow [82] and represent a camera as a bundle of rays. More specifically, each camera is represented by the $16\times 16$ camera rays towards the image patch centers, and each ray by its 6-D Plücker coordinates. Our DPT dense head can be easily adapted to output a $16\times 16$ grid of rays. From the estimated ray bundles, we can then recover camera poses by solving least-square optimizations for camera centers and rotation matrices [82].
Sintel TUM-dyn. ScanNet
DUSt3R [CVPR24] $0.290$ $0.140$ $0.246$
Spann3R [3DV25] $0.329$ $0.056$ $\mathbf{0.096}$
CUT3R [CVPR25] $0.213$ $\mathbf{0.046}$ $0.099$
L4P $\mathbf{0.115}$ $0.098$ $0.150$
For evaluation, we follow [81] and show results on both dynamic (Sintel and TUM-dynamics) and static scenes (ScanNet). We provide ATE (lower is better) and focus on feedforward approaches using general architectures. L4P performs favorably or on par with recent approaches like DUSt3R [65], Spann3R [60], CUT3R [64]. We show visual comparisons in Figure 5 (last row), and discuss additional metrics and methods in the Supplementary.
Discussion. Freezing the video encoder and fine-tuning a task-specific head is a simple way to add a new task. Better strategies may exist that allow for fine-tuning of the video encoder without affecting the performance of other tasks. Note that since we do not estimate in metric units, tasks like camera pose estimation and depth estimation can have different scales, albeit solved within a single unified framework. Adding constraints in the loss functions can help make the outputs of a new task consistent in scale with existing ones. We leave the investigation of these important problems to future work.
@@ 6 Conclusions
We present a unified framework to solve multiple low-level 4D vision perception tasks, both dense and sparse. We achieve this by adopting a strong pre-trained video masked auto-encoder and design lightweight task heads to harness its representation power. Our simple yet versatile designs for task heads allow for effortless and generalizable adaptation to multiple 4D vision perception tasks.
@@ 7 Acknowledgements
In this supplementary document, we provide comparisons for the inference time, and discuss additional implementation details regarding architecture, dataset and training. Please refer to supplementary video for a high-level overview and video results and comparisons.
@@ 1 Inference time
Our method solves multiple perception tasks at once and it is attractive for any application that requires a low-level, holistic understanding of the scene. We show that its performance is as good as, or better than methods designed for individual tasks. Its computational time also makes it appealing, as it solves these tasks jointly, with efficiency comparable with that of single-task methods. For a video clip of size $16\times 224\times 224$, our inference runs in around 300ms on an NVIDIA A6000, which corresponds to $\sim$19ms/frame, or $\sim$28ms/frame if we consider the overhead due to overlapping windows in long videos. This runtime includes all the tasks we discuss in the paper: depth, flow, tracking, motion segmentation, and pose estimation. Table 1 shows the runtime of the SOTA methods we compare against in the paper and confirms that our joint estimation is indeed efficient.
@@ 2 Training datasets
Our video encoder [62] has been pre-trained on 1.35M video clips across various data sources using masked auto-encoding. To fine-tune our model, we use a limited number of synthetic datasets covering a varying range of annotations, and rely on the priors from the video encoder for generalization.
Kubric [16]. This synthetic dataset has multi-object interactions with many random objects. We use it to generate annotations for depth, flow, motion-based segmentation, camera pose estimation and 2D/3D tracking. Each video is 24 frames long, and we use a total of 15k videos from the movi-e and movi-f subsets of the data. Kubric provides meta-data for object and camera trajectories, which could be used to generate 3D tracks. We follow official guidelines and generate the annotations for 3D tracking by sampling around 8-12k tracks in each video.
PointOdyssey [83]. We use this synthetic dataset for depth, camera pose and 2D/3D tracking annotations. The dataset consists of 159 videos, averaging 2k frames long, with human and object interactions under different lighting and atmospheric conditions. We sample smaller video clips from the long videos to form our dataset.
DynamicReplica [26]. We use this synthetic dataset for depth, flow, camera pose and 2D/3D tracking annotations. The dataset consists of 524 videos with humans and animals in motion, and we sample smaller video clips to form our dataset. Since this dataset has higher fps videos, we sample videos with strides of 1, 2 and 4.
TartanAir [68]. Finally, to increase the scene-level data distribution we use TartanAir to generate annotations for flow, depth and camera pose. The data is collected in photo-realistic simulation environments, with both indoor and outdoor scenes, in the presence of various light and weather conditions. We sample smaller video clips from this dataset.
@@
Table 1: Inference time. We compare our per-frame inference time with several task-specific approaches and show at least comparable speed-wise to methods specialized for each task.
@@ 3 Architecture
Our video encoder [62] processes video-clips of size $16\times 224\times 224$. It uses a patch-size of $2\times 14\times 14$, which results in $P=2048$ video tokens, and an embedding dimension of $C=1408$. It has 40 encoder blocks, and we use output from blocks $[14,21,28,36]$ for DPT heads for dense tasks, while our sparse-head uses features from the last block (block 39). Feeding the sparse and dense heads with tokens from different blocks allows us to maintain the performance on dense tasks while we fine-tune our model to train the memory mechanism for the tracking tasks. We adapt the DPT head [42], originally designed to work with image tokens, to work with video tokens. We do this by adding appropriate rearrangement operation that maps three-dimensional video tokens to video feature maps and replace 2D-convolutions with 3D.
For the sparse head, we adapt two-way attention and mask decoding from SAM [29]. We make appropriate modifications to make it work with video tokens and for our point tracking task. For an input point query, we use 3D positional encoding of dimension $C$ to generate a point query token. To identify all our input/output tokens we add learnable embeddings of dimension $C$. These input/output tokens interact with the input video tokens via a two-way attention mechanism. Following SAM, we only keep two layers of two-way attention mechanism. In the mask decoding stage, we add rearrangement operations to map video tokens of size $P\times C$ to video feature maps of size $T\times H\times W\times C/8$. We replace the 2D convolutions in the original mask-decoder of SAM with 3D convolutions. The video feature maps are finally mapped to 2D heatmap, depth-map and visibility map by doing a dot-product with the corresponding output tokens processed via the two-stage attention mechanism, as shown in Figure 3 in the main paper.
To enable memory mechanism for online tracking using overlapping windows, we introduce a track-feature token $\mathcal{F}$ that is passed to subsequent windows as an additional prompt. We do this by mapping the processed feature token $\mathcal{F}$ from the previous window via a linear layer with input and output dimensions of $C$ and then passing it as input to the next window. To provide even more cross-window information, we pass the video tokens decoded by the two-way attention stage of the current window to the next. We achieve this by projecting the decoded video tokens in the overlapping region via a linear layer with input and output dimensions of $C$, and by adding them to the corresponding incoming video tokens to the two-way attention stage in the next window. For the non-overlapping region we add a learnable mask token of dimension $C$.
@@ 4 Training
We initialize our video encoder using a pretrained VideoMAE [62] and fine-tune our model in three stages. All the stages use a batch size of 8 and are trained on a single 8-GPU (NVIDIA A100) node. In the first stage we train all the parameters of our model for depth, flow and 3D tracking on a single window of $T=16$ frames and train only on the Kubric dataset. In the second stage of training, we add the remaining three datasets and further fine-tune the model on a single window for all the tasks. In this stage, since not all the datasets include all the tasks (flow missing in PointOdyssey and point tracking missing in TartanAir), we formed our batch by sampling two videos from each of the four datasets. This ensures a fixed batch-size per-task across training iterations. In the first two stages, for each video we estimate all $T$ frames for dense tasks and for 3D tracking we construct a batch of 80 tracks and select point queries randomly across the visible parts of tracks.
In the third stage, we further fine-tune our model for the tracking tasks using unrolled-window training and the memory mechanism for online tracking. We train on videos of length $S=40$, windows of size 16 frames and a stride of 8, which results in 4 overlapping windows, and we construct a batch of 48 tracks during this stage, but generate point queries only in the first 20 frames to force the network to learn long-range tracking. To train this stage efficiently, we only train for the 2D and 3D tracking tasks and freeze all the parameters, except the last three layers (37-39) of the video encoder and the sparse head. Note that since other dense tasks use the output of layers smaller than 37, their performance remains unaffected by this stage. For all the stages we use AdamW optimizer with a maximum learning rate of $5\times 10^{-5}$, beta values of $(0.9,0.999)$, weight-decay value of $0.05$, gradient-clipping value of $0.05$ and trained using mixed-precision. We use cosine annealing strategy for learning rate where we use first 5k iterations to increase the learning rate to the maximum value. Training takes around four days for all the stages, one day each for the first two stages and two days for the third stage.
@@
5 Motion-based segmentation
For motion-based segmentation, we generate GT annotations for training using Kubric dataset. We use the camera-pose and the 3D track information to detect which 3D tracks come from dynamic vs. static objects. We combine this information with the provided rigid-object segmentations to generate annotations for motion-based segmentation. For evaluation datasets, we use Virtual KITTI [5] and Spring [37] datasets. Both datasets provide camera poses, depth and scene-flow information. We combine depth and camera information to estimate scene-flow due to the camera-motion. This allows us to compute scene-flow due to dynamic objects and generate ground-truth annotations for motion-based segmentation. We train the DPT head for this task using a binary cross-entropy loss. We use a batch size of 8, use random affine augmentations, and train for 25k iterations.
6 Camera pose estimation
Here we provide additional details and results for the pose estimation task.
Sintel ($\sim$50 frames) TUM-dynamics (90 frames) ScanNet (90 frames)
Method ATE $\downarrow$ RPE-trans $\downarrow$ RPE-rot $\downarrow$ ATE $\downarrow$ RPE-trans $\downarrow$ RPE-rot $\downarrow$ ATE $\downarrow$ RPE-trans $\downarrow$ RPE-rot $\downarrow$
DROID-SLAM∗† [56] $0.175$ $0.084$ $1.912$ $\mathbf{0.021}$ $\mathbf{0.013}$ $\mathbf{0.358}$ $\mathbf{0.046}$ $\mathbf{0.015}$ $\mathbf{0.449}$
DUSt3R [65] $0.417$ $0.250$ $5.796$ $0.083$ $0.017$ $3.567$ $0.081$ $0.028$ $0.784$
MonST3R [81] $\mathbf{0.111}$ $\mathbf{0.044}$ $\mathbf{0.869}$ $0.098$ $0.019$ $0.935$ $0.077$ $0.018$ $0.529$
DUSt3R [65] (w/o GA) $0.290$ $0.132$ $7.869$ $0.140$ $0.106$ $3.286$ $0.246$ $0.108$ $8.210$
Spann3R [60] $0.329$ $0.110$ $4.471$ $0.056$ $0.021$ $0.591$ $\mathbf{0.096}$ $0.023$ $0.661$
CUT3R [64] $0.213$ $0.066$ $\mathbf{0.621}$ $\mathbf{0.046}$ $\mathbf{0.015}$ $\mathbf{0.473}$ $0.099$ $\mathbf{0.022}$ $\mathbf{0.600}$
L4P $\mathbf{0.115}$ $0.061$ $2.491$ $0.098$ $0.017$ $3.296$ $0.150$ $0.036$ $3.862$
L4P w/ intrinsics $\mathbf{0.115}$ $\mathbf{0.039}$ $0.844$ $0.099$ $\mathbf{0.015}$ $1.047$ $0.150$ $0.031$ $1.093$
Table 2: Additional camera pose estimation results. We consider two types of approaches: optimization-based and feedforward approaches. $\dagger$ indicates methods that require camera intrinsics at inference time, and * indicates the use of global optimization. “DUSt3R (w/o GA)” is an online variant of DUSt3R, which aligns all video frames with first frame, without using global alignment (GA).
Architecture. We modify the DPT head used in the previous dense tasks to output $16\times 16$ instead of full dense resolution, matching the resolution suggested in [82]. This is achieved by changing the sampling rates at several layers within the DPT head, The last layer produces 6 channels, representing the 6-D Plücker coordinates $(d,o\times d)$, where $d$ is the ray direction and $o$ is a point on the ray.
Training. We freeze the video encoder and train the camera pose estimation head with all datasets described in Section 2. GT rays are constructed for all samples using provided camera intrinsics and poses. We apply L2-loss on the Plücker coordinates, and use 16-frame clips for training. We use a batch size of 8 and train for 250k iterations.
Testing. At inference time, output Plücker coordinates can be converted to traditional camera pose representations for evaluation by solving least-square optimizations. Following [82], we consider and evaluate under two settings, one where camera intrinsics are provided at inference time, and one where they are not. The evaluation datasets, Sintel, TUM-dynamics, and ScanNet, all contain sequences longer than 16 frames, and therefore require inference beyond a single $T=16$ window. In order to do so, we perform inference on 16-frame windows with stride 8, and align the predictions using similarity transformation between consecutive windows. The similarity transformation is estimated from the overlapping frames to minimize alignment errors. We compare with both optimization-based and feedforward approaches in Table 2.